# Hume: Introducing System-2 Thinking in Visual-Language-Action Model

### v2

@@ 1 Introduction
Creating generalist robots to perform various tasks like humans in the physical world has been a long-standing goal (livision, ; brohan2023rt, ; kim2024openvla, ; li2024cogact, ; zheng2024tracevla, ; li2024towards, ; zheng2025universal, ; pertsch2025fast, ; geminiroboticsteam2025geminiroboticsbringingai, ; yao2025thinksmallactbig, ). Cognitive psychologists have revealed that humans conduct a deep, deliberate form of thinking when tackling complex problems, such as mathematical proofs or dinner making. This slow and reflective thought process is known as System-2 thinking, while the fast thinking process that relies on intuition is called System-1 thinking kahneman2011thinking . Inspired by the dual process theory of human cognition, thinking and reasoning steps have been introduced to enhance LLMs’ capability to solve complex problems in digital domains and achieved significant results. Intuitively, generalist robots in the physical world also require similar System-2 slow thinking capabilities to perform dynamic, dexterous, and complex tasks, while intuition-based fast System-1 thinking is not capable of tackling delicate, tenuous, and fallible robot action prediction. Therefore, a crucial question for generalist robot policies designed to solve complex robotic tasks in diverse scenarios is: how to enable effective System-2 slow thinking in generalist robot policy for accurate action prediction?
@@
In this work, we introduce a dual-system Vision-Language-Action model Hume, which powers the VLA model with System-2 thinking capabilities through value-guided repeat sampling and cascaded action denoising. The System 2 of Hume is built on top of a pre-trained vision-language model (VLM) and attach it with two specialized heads: a flow matching denoising head to predict robot actions, and a value-query head to estimate the state-action value of predicted actions. It processes the robot’s observation and language instruction to predict long-horizon action chunk through action denoising head. Subsequently, the corresponding state-action value are estimated conditioning on predicted action chunk. The value-guided thinking is conducted by repeat sampling multiple action chunks and selecting one with the highest value. For System 1 of Hume, it takes one short segment from the selected long-horizon action chunks from System 2, current visual observation, and robot states, then conducts cascaded action denoising to generate final fluid robot actions via a separate lightweight diffusion policy. At deployment time, System 2 performs value-guided thinking at a low frequency (4 Hz) while System 1 asynchronously receives the System 2 selected action chunk and predicts fluid actions in real time (90 Hz). Equipped with the proposed value-guided thinking and cascaded action denoising, Hume explores powerful System 2 slow thinking paradigm to enhance the VLA models. We extensively evaluate and ablate Hume on both standard simulation benchmarks and real-robot platforms, including 21 real-world robot settings and 3 simulation environments. To validate Hume’s capability to solve complex robot control tasks with the assistance of System-2 Thinking, the test scenarios include variations in viewpoint, texture, lighting, layout, unseen objects, unseen environments, as well as the most challenging humanoid robot control tasks. In summary, the main contributions of this work are three-folds:
@@ 2 Related Work
Dual-System Vision-Language-Action Models. Recently, several studies (livision, ; brohan2023rt, ; kim2024openvla, ; li2024cogact, ; zheng2024tracevla, ; li2024towards, ; zheng2025universal, ; pertsch2025fast, ; bu2024towards, ; helix, ; han2024dual, ) have extended VLMs for robot control. RT-2 (brohan2023rt, ) fine-tunes PaLI-X (chen2023pali, ) with discretized action tokens, while OpenVLA (kim2024openvla, ) adapts Prismatic VLM (karamcheti2024prismatic, ) on the OXE dataset (o2024open, ). $\pi_{0}$ (black2024pi_0, ) integrates PaliGemma with flow-matching for continuous actions. To address efficiency and integration challenges, dual-system architectures have emerged. HiRT (zhang2024hirt, ) runs VLMs at low frequencies while maintaining high-frequency vision-based control for real-time interaction. DexVLA (wen2025dexvla, ) uses diffusion action experts with embodiment curriculum learning across multiple robot types. GR00T N1 (bjorck2025gr00t, ) features an end-to-end trained dual-system specifically for humanoid robots. Gemini Robotics (geminiroboticsteam2025geminiroboticsbringingai, ) builds on Gemini 2.0 with specialized models for control and reasoning. HiRobot (shi2025hirobotopenendedinstruction, ) enables processing of complex instructions with situated feedback. These approaches improve efficiency, success rates, and adaptability compared to monolithic architectures.
@@
System 2 and System-2 Thinking. Numerous studies (wei2022chain, ; yao2023treethoughtsdeliberateproblem, ; shinn2023reflexionlanguageagentsverbal, ; hao2023reasoninglanguagemodelplanning, ; shao2024deepseekmathpushinglimitsmathematical, ; miao2023selfcheck, ; Tree_of_Thought, ; liu2024groupdebateenhancingefficiencymultiagent, ; song2025progcoprogramhelpsselfcorrection, ) have explored System 2 reasoning approaches to enhance LLMs’ problem-solving capabilities. Chain-of-Thought (wei2022chain, ) introduces intermediate reasoning steps before producing answers, while Tree-of-Thoughts (Tree_of_Thought, ) explores multiple solution paths with self-verification. Reflexion (shinn2023reflexionlanguageagentsverbal, ) enables verbal reflection on previous attempts. System-2 Thinking frameworks explicitly model human-like deliberative processes. SETS (chen2025setsleveragingselfverificationselfcorrection, ) combines self-critique with multiple reasoning paths and majority voting. SC-MCTS (gao2024interpretablecontrastivemontecarlo, ) combines multiple reward models for more robust tree search. Addressing efficiency concerns, O1-Pruner (luo2025o1prunerlengthharmonizingfinetuningo1like, ) introduces length-penalty loss to create concise reasoning processes. While these approaches have improved reasoning in language tasks, their application to visual-language-action models remains largely unexplored.
@@
Cascaded Denoising. Cascaded Denoising ho2021cascaded was first proposed as a diffusion model that generates higher-resolution images through a cascading approach. Starting with a diffusion model at low resolution, it continuously upsamples the generated images to obtain high-resolution results. f-DM gu2022fdm applies inter-stage transformations for progressive signal restoration through function learning. Cas-DM an2024bring cascades noise-prediction and image-prediction modules to integrate perceptual losses in diffusion models. SCDM chen2024spectral cascades generation across spectral dimensions, reconstructing hyperspectral bands progressively. HiFI hur2025high cascades consistent-resolution patches for memory-efficient high-resolution frame interpolation. CDM-VTON li2025cascaded employs two-stage cascading for virtual try-on applications. While these approaches have advanced capabilities in image domains, the application of cascaded denoising to integrate System 1 and System 2 remains largely unexplored.
@@ B.1 Value Objective Functions
Formally, the goal of the value-query Head is to learn $Q_{\theta}(\mathbf{q}_{t},\mathbf{A}_{t})$, which is the optimal estimate of the state-action value function ${Q^{\pi}(\mathbf{q}_{t},\mathbf{A}_{t})={\frac{1}{1-\gamma}\sum_{t}\mathbb{E}_{\mathbf{A}_{t}\sim\pi(\mathbf{q}_{t})}\left[\gamma^{t}r(\mathbf{q}_{t},\mathbf{A}_{t})|\mathbf{q}_{t_{0}}=\mathbf{q}_{t},\mathbf{A}_{t_{0}}=\mathbf{A}_{t}\right]}}$ in a Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\rho,\gamma)$. Here $\mathcal{S},\mathcal{A}$ denote the state and action spaces, while $P(\mathbf{q}^{\prime}|\mathbf{q},\mathbf{A})$ and $r(\mathbf{q}_{t},\mathbf{A}_{t})$ are the dynamics and reward functions. $\rho(\mathbf{q})$ denotes the initial state distribution, and $\gamma\in(0,1)$ denotes the discount factor. The training objective of the Value-Query Head is to minimize the Bellman error with a regularization term $\mathcal{R}(\theta)$, which is defined as:

### v4

@@ 1 Introduction
Creating generalist robots to perform various tasks like humans in the physical world has been a long-standing goal (livision, ; brohan2023rt, ; liu2024fastumi, ; kim2024openvla, ; gao2025learning, ; pertsch2025fast, ; yao2024improving, ; zhang2025moma, ; gao2023orla, ; zhang2024decentralized, ; geminiroboticsteam2025geminiroboticsbringingai, ; black2024pi_0, ). Cognitive psychologists have revealed that humans conduct a deep, deliberate form of thinking when tackling complex problems, such as mathematical proofs or dinner making. This slow and reflective thought process is known as System-2 thinking, while the fast thinking process that relies on intuition is called System-1 thinking kahneman2011thinking . Inspired by the dual process theory of human cognition, thinking and reasoning steps have been introduced to enhance LLMs’ capability to solve complex problems in digital domains and achieved significant results. Intuitively, generalist robots in the physical world also require similar System-2 slow thinking capabilities to perform dynamic, dexterous, and complex tasks (liu2024coherent, ; xia2024kinematic, ; wang2025more, ; wang2025skillnavenhancednavigationversatile, ), while intuition-based fast System-1 thinking is not capable of tackling delicate, tenuous, and fallible robot action prediction (xia2025roboticpolicylearninghumanassisted, ; xia2025phoenix, ; yao2025think, ; Lv_2025_CVPR, ). Therefore, a crucial question for generalist robot policies designed to solve complex robotic tasks in diverse scenarios is: how to enable effective System-2 slow thinking in generalist robot policy for accurate action prediction?
@@
In this work, we introduce a dual-system Vision-Language-Action model Hume, which powers the VLA model with System-2 thinking capabilities through value-guided repeat sampling and cascaded action denoising. The System 2 of Hume is built on top of a pre-trained vision-language model (VLM) and attach it with two specialized heads: a flow matching lipman2022flow ; lin2025beyond denoising head to predict robot actions, and a value-query head to estimate the state-action value of predicted actions. It processes the robot’s observation and language instruction to predict long-horizon action chunk through action denoising head. Subsequently, the corresponding state-action value are estimated conditioning on predicted action chunk. The value-guided thinking is conducted by repeat sampling multiple action chunks and selecting one with the highest value. For System 1 of Hume, it takes one short segment from the selected long-horizon action chunks from System 2, current visual observation, and robot states, then conducts cascaded action denoising to generate final fluid robot actions via a separate lightweight diffusion policy. At deployment time, System 2 performs value-guided thinking at a low frequency (4 Hz) while System 1 asynchronously receives the System 2 selected action chunk and predicts fluid actions in real time (90 Hz). Equipped with the proposed value-guided thinking and cascaded action denoising, Hume explores powerful System 2 slow thinking paradigm to enhance the VLA models. We extensively evaluate and ablate Hume on both standard simulation benchmarks and real-robot platforms, including 21 real-world robot settings and 3 simulation environments. To validate Hume’s capability to solve complex robot control tasks with the assistance of System-2 Thinking, the test scenarios include variations in viewpoint, texture, lighting, layout, unseen objects, unseen environments, as well as the most challenging humanoid robot control tasks. In summary, the main contributions of this work are three-folds:
@@ 2 Related Work
Dual-System Vision-Language-Action Models. Recently, several studies (livision, ; brohan2023rt, ; kim2024openvla, ; li2024cogact, ; zeng2025FSDrive, ; zheng2024tracevla, ; pmlr-v235-lv24a, ; li2024towards, ; zheng2025universal, ; liu2023learning, ; pertsch2025fast, ; bu2024towards, ; helix, ; han2024dual, ) have extended VLMs for robot control. RT-2 (brohan2023rt, ) fine-tunes PaLI-X (chen2023pali, ) with discretized action tokens, while OpenVLA (kim2024openvla, ) adapts Prismatic VLM (karamcheti2024prismatic, ) on the OXE dataset (o2024open, ). $\pi_{0}$ (black2024pi_0, ) integrates PaliGemma with flow-matching for continuous actions. To address efficiency and integration challenges, dual-system architectures have emerged. HiRT (zhang2024hirt, ) runs VLMs at low frequencies while maintaining high-frequency vision-based control for real-time interaction. DexVLA (wen2025dexvla, ) uses diffusion action experts with embodiment curriculum learning across multiple robot types. GR00T N1 (bjorck2025gr00t, ) features an end-to-end trained dual-system specifically for humanoid robots. Gemini Robotics (geminiroboticsteam2025geminiroboticsbringingai, ) builds on Gemini 2.0 with specialized models for control and reasoning. HiRobot (shi2025hirobotopenendedinstruction, ) enables processing of complex instructions with situated feedback. These approaches improve efficiency, success rates, and adaptability compared to monolithic architectures.
@@
System 2 and System-2 Thinking. Numerous studies (wei2022chain, ; shinn2023reflexionlanguageagentsverbal, ; hao2023reasoninglanguagemodelplanning, ; shao2024deepseekmathpushinglimitsmathematical, ; zhang2025cross, ; NEURIPS2024_0d9dcd4e, ; miao2023selfcheck, ; wang2025morphmark, ; Tree_of_Thought, ) have explored System 2 reasoning approaches to enhance LLMs’ problem-solving capabilities. Chain-of-Thought (wei2022chain, ; zhaxizhuoma2024alignbot, ) introduces intermediate reasoning steps before producing answers, while Tree-of-Thoughts (Tree_of_Thought, ) explores multiple solution paths with self-verification. Reflexion (shinn2023reflexionlanguageagentsverbal, ) enables verbal reflection on previous attempts. System-2 Thinking frameworks explicitly model human-like deliberative processes. SETS (chen2025setsleveragingselfverificationselfcorrection, ) combines self-critique with multiple reasoning paths and majority voting. SC-MCTS (gao2024interpretablecontrastivemontecarlo, ) combines multiple reward models for more robust tree search. Addressing efficiency concerns, O1-Pruner (luo2025o1prunerlengthharmonizingfinetuningo1like, ) introduces length-penalty loss to create concise reasoning processes. While these approaches have improved reasoning in language tasks, their application to visual-language-action models remains largely unexplored.
@@
Cascaded Denoising. Cascaded Denoising ho2021cascaded was first proposed as a diffusion model zhang2025revisiting that generates higher-resolution images through a cascading approach. Starting with a diffusion model at low resolution, it continuously upsamples the generated images to obtain high-resolution results. f-DM gu2022fdm applies inter-stage transformations for progressive signal restoration through function learning. Cas-DM an2024bring cascades noise-prediction and image-prediction modules to integrate perceptual losses in diffusion models. SCDM chen2024spectral cascades generation across spectral dimensions, reconstructing hyperspectral bands progressively. HiFI hur2025high cascades consistent-resolution patches for memory-efficient high-resolution frame interpolation. CDM-VTON li2025cascaded employs two-stage cascading for virtual try-on applications. While these approaches have advanced capabilities in image domains, the application of cascaded denoising to integrate System 1 and System 2 remains largely unexplored.
@@ B.1 Value Objective Functions
Formally, the goal of the value-query Head is to learn $Q_{\theta}(\mathbf{q}_{t},\mathbf{A}_{t})$, which is the optimal estimate (liu2023improving, ) of the state-action value function ${Q^{\pi}(\mathbf{q}_{t},\mathbf{A}_{t})={\frac{1}{1-\gamma}\sum_{t}\mathbb{E}_{\mathbf{A}_{t}\sim\pi(\mathbf{q}_{t})}\left[\gamma^{t}r(\mathbf{q}_{t},\mathbf{A}_{t})|\mathbf{q}_{t_{0}}=\mathbf{q}_{t},\mathbf{A}_{t_{0}}=\mathbf{A}_{t}\right]}}$ in a Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\rho,\gamma)$. Here $\mathcal{S},\mathcal{A}$ denote the state and action spaces, while $P(\mathbf{q}^{\prime}|\mathbf{q},\mathbf{A})$ and $r(\mathbf{q}_{t},\mathbf{A}_{t})$ are the dynamics and reward functions. $\rho(\mathbf{q})$ denotes the initial state distribution, and $\gamma\in(0,1)$ denotes the discount factor. The training objective of the Value-Query Head is to minimize the Bellman error with a regularization term $\mathcal{R}(\theta)$, which is defined as:
@@ C.1 Simulation Benchmark Details
Google Robot Evaluation Setup Policy Pick Coke Can Move Near Open / Close Drawer #Overall
Horizontal
Laying
Vertical
Laying
Standing
Average
Average
Open
Close
Average
Average